OpenStudy (anonymous):

A distribution center receives shipments from three different factories in quantities of 50, 35, and 20. Three times a product is selected at random. Find the probability that all three products came from the second factory.

OpenStudy (anonymous):

1/3 times1/3times1/3 = 1/27

OpenStudy (anonymous):

oh i don't think so

OpenStudy (anonymous):

ok now i think so. sorry

OpenStudy (anonymous):

haha i didnt think so

OpenStudy (anonymous):

Total number is 50+35+20=105. I believe it should be \[(\frac{35}{105})^3\]

OpenStudy (anonymous):

Oh, whoops, that's the same thing that was written before.

OpenStudy (anonymous):

haha my problem changes everytime i answer it wrong. now the numbers are 50, 40, and 25

OpenStudy (anonymous):

and it wants the probability that they came from the third factory

OpenStudy (anonymous):

Well, the way I did it should work. Find the sum of all the units, then find the probability that a single unit came from the 3rd factory, and then cube the probability.

OpenStudy (anonymous):

It said the answer was 0.009 is that what you woulda got?

OpenStudy (anonymous):

I got 0.010

OpenStudy (anonymous):

yeah i didnt get the answer they wanted either haha

OpenStudy (anonymous):

Hm...I wonder why they're getting that.

OpenStudy (anonymous):

It was a weird process but I figured it out, thanks for trying